a- Combustion of 13.5 g of sample of an hydrocarbon yields 24.41 g CO, and 14.49 g H2O. The molar mass of the compound is 246 g/mol. What are the empirical and molecular formulas?

b- Balance the combustion equation using the molecular formula.


1
Expert's answer
2021-05-11T05:59:35-0400

"a) \ number \ of\ moles \ of \ carbon=\dfrac{24.41}{28}=0.8714 \ moles"

number of moles of hydrogen = "\dfrac{14.49}{18}\times 2=0.805\times2=1.61 \ moles"


To get the whole number of mole we divide it by - 0.8714


we get moles of carbon = 1


we get moles of hydrogen = 1.84 = 2


multiplying by factor 2 to get the whole number-


we get empirical formula as = "C_2H_4"


Molecular formula = to find molecular formula , we determine how many empirical formula it takes to make 1 molecular formula =


molar mass of "C_2H_4= 28 \ g/ mol"


"=\dfrac{246}{28}=""\ 8.78"


molecular fomula will be "C_{16}H_{32}"


B) To balance the combustion equation using molecular formula-


"C_{16}H_{32} \ + 32\ O\to" "16 \ CO+ 16\ H_2O"